Korean22:

so im suppose to find the slop of y=-7x +11

surjithayer:

y=mx+c where m is the slope.

XLynx:

-7. Slope-intercept is written in y = mx+b or in special cases, f(x) = mx+b. Make sure to carry the negative. :)

Korean22:

tysm ^^

XLynx:

f(x) is written if it's a function. If there are several Y values for one X, it isn't a function unless you restrict it. A good way to check is the vertical line test. If it hits more than one point on the graph, it isn't a function. Just draw a line through the function. :)

Korean22:

oh okay so the funtion is -7?

XLynx:

In a way. You have to have the Y intercept as well as the X. f(x) = -7 isn't a function since literally everything is on the vertical line XD It's f(x) = -7x+11.

Korean22:

wow ur smart

XLynx:

The slope is -7, the y intercept is 11.

XLynx:

No, I'm not. I'm a total idiot. XD

Korean22:

no ur not

XLynx:

I'm not smart in any fashion. What's your rhyme behind your reason?

Korean22:

hmm